# DESY-24-134 TIF-UNIMI-2024-17 Edinburgh 2024/9

### v1

@@ Acknowledgments
S. F. performed this work in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant PHY-2210452, and is partially supported by an Italian PRIN2022 grant. R. D. B, L. D. D. and R. S. thank STFC for support via grant awards ST/T000600/1 and ST/X000494/1. L. H.-L. and R. S. T. thank STFC for support via grant awards ST/T000856/1 and ST/X000516/1. T. C. acknowledges that this project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). E. R. N. is supported by the Italian Ministry of University and Research (MUR) through the “Rita Levi-Montalcini” Program. F. H. is supported by the Academy of Finland project 358090 and is funded as a part of the Center of Excellence in Quark Matter of the Academy of Finland, project 346326. M.U. is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation Programme (grant agreement n.950246), and partially by the STFC consolidated grant ST/T000694/1 and ST/X000664/1.

### v2

@@ Acknowledgments
S. F. performed this work in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant PHY-2210452, and is partially supported by an Italian PRIN2022 grant. R. D. B, L. D. D. and R. S. thank the Science and Technology Facilities Council (STFC) for support via grant awards ST/T000600/1 and ST/X000494/1. L. H.-L. and R. S. T. thank the Science and Technology Facilities Council (STFC) for support via grant awards ST/T000856/1 and ST/X000516/1. T. C. acknowledges that this project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). E. R. N. is supported by the Italian Ministry of University and Research (MUR) through the “Rita Levi-Montalcini” Program. F. H. is supported by the Academy of Finland project 358090 and is funded as a part of the Center of Excellence in Quark Matter of the Academy of Finland, project 346326. M.U. is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation Programme (grant agreement n.950246), and partially by the STFC consolidated grant ST/T000694/1 and ST/X000664/1.